# FIGRET: Fine-Grained Robustness-Enhanced Traffic Engineering

### v1

@@ D.3. Drawbacks of Using RL
Network Precomp. time (s)
FIGRET TEAL OBL COPE
GEANT (#Nodes 23) 150 1500 100 1000
ToR DB (#Node 155) 500 7000 - -
ToR WEB (#Node 324) 1500 20000 - -
Table 8. Comparing the precomputation time, OBL represents Oblivious. A ‘-’ indicates that the corresponding TE scheme could not be solved within the constraints of the maximum memory of the given machine and a solving time limit of one day.
@@ D.4. Implementing FIGRET
DNN architecture. Similar to DOTE, except for the input and output layers, FIGRET uses five fully connected neural network layers with 128 neurons each and employs ReLU(x) activation, except for the output layer, which uses Sigmoid(x). Optimizer. During the training process, FIGRET employs the Adam optimizer (Kingma and Ba, 2014) for stochastic gradient descent.
@@ Appendix E Additional link failure results
In this section, we present the results from evaluating link failures on pFabric and ToR-level DC Facebook DB, as shown in Figures 12 and 13. As can be seen in Figure 13, when the network traffic demands exhibit high dynamics, the current TE scheme in Google’s Jupiter, even with the knowledge of which links are going to fail, fails to achieve satisfactory results.
@@
Figure 13. Coping with different numbers of random link failures on ToR-level Facebook DB.
@@ Appendix F Visualization of traffic demands
To visually demonstrate the extent of traffic data migration over time, we conduct a visualization analysis of traffic demands from both the PoD-level Facebook DB and the ToR-level Facebook DB using the t-distributed stochastic neighbor embedding (t-SNE) method (Hinton and Roweis, 2002) in this section. We analyze their traffic from $0\%$ to $100\%$ using t-SNE and compare the differences in the intervals of $0\%-25\%$, $25\%-50\%$, $50\%-75\%$, and $75\%-100\%$. The 2-dimensional t-SNE components are plotted in Figure 14 and Figure 15.
@@
(a) 0%-25%
(b) 25%-50%
(c) 50%-75%
(d) 75%-100%
Figure 14. Visualizing Traffic Demands of the PoD-level Facebook DB using the t-SNE Method.
(a) 0%-25%
(b) 25%-50%
(c) 50%-75%
(d) 75%-100%
Figure 15. Visualizing Traffic Demands of the ToR-level Facebook DB using the t-SNE Method.
@@ G.1. Objective mismatch in Demand-prediction-based TE
In demand-prediction-based TE, a notable mismatch exists between the objective of accurately predicting future traffic and minimizing the MLU (Perry et al., 2023). This mismatch primarily arises from the influence of network topology on MLU, indicating that traffic demands exert varied impacts on MLU. For instance, traffic traversing paths with higher capacity tends to have a lesser effect on MLU. Similarly, traffic flowing through paths without shared edges with other paths will also impact MLU to a lesser degree. An example illustrating this concept is provided in Figure 16.
(a) In the network topology, where numbers on edges indicate link capacity, traffic demands of $d_{1}$ and $d_{2}$ need to be sent from $s$ to $t_{1}$ and $s$ to $t_{2}$, respectively.
(b) An example concerning predicted traffic, where the split ratios are only specified for the paths $s\rightarrow t_{1}$ and $s\rightarrow t_{2}$, with the split ratios for the other two paths being $1$ minus the corresponding split ratios.
Figure 16. An example demonstrating the mismatch between achieving accurate traffic predictions and minimizing MLU. Two predicted demands achieve the same level of prediction accuracy (with identical mean-square-error values). However, the optimal split ratios derived from these two traffic predictions lead to different MLUs. Incorrect predictions of the traffic from $s$ to $t_{2}$ have a lesser impact on MLU. This is due to the larger capacity of the path from $s$ to $t_{2}$.
@@ G.2. Limited gains from window expansion
(a) H = 12
(b) H = 64
Figure 17. Cosine similarity analysis using a window of H historical TMs vs. the currently-seen TM.
Deep learning approaches demonstrate powerful capabilities in using information within a window to obtain a better $D^{\text{expect}}$. Therefore, in this section, we consider whether, under the premise of sufficient computational power and memory, expanding the window size could make $D^{\text{expect}}$ highly accurate and prevent mis-predictions, thereby eliminating the need to consider sudden traffic changes. Using an analysis method similar to §5, we examine the cosine similarity between upcoming traffic and traffic from the past $H$ time windows, where $H$ is increased to 64. The results are summarized in Figure 17. However, the results indicate that, regardless of whether in stable or bursty network types, the cosine similarity does not significantly increase compared to when $H$ is 12, suggesting that expanding the window size is ineffective in avoiding sudden traffic changes.
Due to the limited information within the window, DOTE, which solely relies on window-based information and prioritizes MLU, may result in suboptimal configurations. An example is shown in Figure 18, which illustrates an inadequate network configuration by DOTE on the ToR-level Facebook DB (with an Omniscient-normalized MLU of 2.86). We selected the SD pair responsible for the MLU, and as seen in Figure 18(a), the traffic on this SD pair was consistently stable at a relatively low value. Therefore, DOTE assigned a configuration for this SD pair involving paths with high path sensitivity. However, in the next time snapshot, the SD pair burst unexpectedly, significantly impacting the highly sensitive paths, leading to high MLU.
(a) The arrival scenario of the traffic for a SD pair.
(b) Path sensitivity of this SD pair.
Figure 18. Showcasing the limitations of the DOTE approach.

### v3

@@ D.4. Implementing FIGRET
DNN architecture. Similar to DOTE, except for the input and output layers, FIGRET uses five fully connected neural network layers with 128 neurons each and employs ReLU(x) activation, except for the output layer, which uses Sigmoid(x).
Optimizer. During the training process, FIGRET employs the Adam optimizer (Kingma and Ba, 2014) for stochastic gradient descent.
@@ Appendix E Additional link failure results
In this section, we present the results from evaluating link failures on pFabric and ToR-level DC Meta DB, as shown in Figures 23 and 24. As can be seen in Figure 24, when the network traffic demands exhibit high dynamics, the current TE scheme in Google’s Jupiter data centers, even with the knowledge of which links are going to fail, fails to achieve satisfactory results.
@@
Figure 24. Coping with different numbers of random link failures on ToR-level Meta DB.
@@ Appendix F Visualization of traffic demands
To visually demonstrate the extent of traffic data migration over time, we conduct a visualization analysis of traffic demands from both the PoD-level Meta DB and the ToR-level Meta DB using the t-distributed stochastic neighbor embedding (t-SNE) method (Hinton and Roweis, 2002) in this section. We analyze their traffic from $0\%$ to $100\%$ using t-SNE and compare the differences in the intervals of $0\%-25\%$, $25\%-50\%$, $50\%-75\%$, and $75\%-100\%$. The 2-dimensional t-SNE components are plotted in Figure LABEL:fig:_t-sne_facebook_pod_a and Figure LABEL:fig:_t-sne_facebook_tor_a.
@@ G.1. Objective mismatch in Demand-prediction-based TE
In demand-prediction-based TE, a notable mismatch exists between the objective of accurately predicting future traffic and minimizing the MLU (Perry et al., 2023). This mismatch primarily arises from the influence of network topology on MLU, indicating that traffic demands exert varied impacts on MLU. For instance, traffic traversing paths with higher capacity tends to have a lesser effect on MLU. Similarly, traffic flowing through paths without shared edges with other paths will also impact MLU to a lesser degree. An example illustrating this concept is provided in Figure LABEL:fig:_mismatch_in_prediction.
@@ G.2. Limited gains from window expansion
Deep learning approaches demonstrate powerful capabilities in using information within a window to obtain a better $D^{\text{expect}}$. Therefore, in this section, we consider whether, under the premise of sufficient computational power and memory, expanding the window size could make $D^{\text{expect}}$ highly accurate and prevent mis-predictions, thereby eliminating the need to consider sudden traffic changes. Using an analysis method similar to §5, we examine the cosine similarity between upcoming traffic and traffic from the past $H$ time windows, where $H$ is increased to 64. The results are summarized in Figure LABEL:fig:_window_of_H. However, the results indicate that, regardless of whether in stable or bursty network types, the cosine similarity does not significantly increase compared to when $H$ is 12, suggesting that expanding the window size is ineffective in avoiding sudden traffic changes.
Due to the limited information within the window, DOTE, which solely relies on window-based information and prioritizes MLU, may result in suboptimal configurations. An example is shown in Figure LABEL:fig:_showcasing_the_limitations_of_DOTE, which illustrates an inadequate network configuration by DOTE on the ToR-level Meta DB (with an Omniscient-normalized MLU of 2.86). We selected the SD pair responsible for the MLU, and as seen in Figure LABEL:fig:_dote_arrival_scenario, the traffic on this SD pair was consistently stable at a relatively low value. Therefore, DOTE assigned a configuration for this SD pair involving paths with high path sensitivity. However, in the next time snapshot, the SD pair burst unexpectedly, significantly impacting the highly sensitive paths, leading to high MLU.